### v3

# DrivAer Transformer: A high-precision and fast prediction method for vehicle aerodynamic drag coefficient based on the DrivAerNet++ dataset

Abstract
At the current stage, deep learning-based methods have demonstrated excellent capabilities in evaluating aerodynamic performance, significantly reducing the time and cost required for traditional computational fluid dynamics (CFD) simulations. However, when faced with the task of processing extremely complex three-dimensional (3D) vehicle models, the lack of large-scale datasets and training resources, coupled with the inherent diversity and complexity of the geometry of different vehicle models, means that the prediction accuracy and versatility of these networks are still not up to the level required for current production. In view of the remarkable success of Transformer models in the field of natural language processing and their strong potential in the field of image processing, this study innovatively proposes a point cloud learning framework called DrivAer Transformer (DAT). The DAT structure uses the DrivAerNet++ dataset, which contains high-fidelity CFD data of industrial-standard 3D vehicle shapes. enabling accurate estimation of air drag directly from 3D meshes, thus avoiding the limitations of traditional methods such as 2D image rendering or signed distance fields (SDF). DAT enables fast and accurate drag prediction, driving the evolution of the aerodynamic evaluation process and laying the critical foundation for introducing a data-driven approach to automotive design. The framework is expected to accelerate the vehicle design process and improve development efficiency.
preprint: AIP/123-QED
I Introduction
With the rapid development of electric vehicles in China, automakers are launching new models faster and faster, and the vehicle development cycle is shortened, which puts forward higher requirements for efficient and accurate aerodynamic calculations. However, traditional automotive aerodynamics relies on the combination of computational fluid dynamics (CFD) and wind tunnel tests, which is a less efficient method and cannot meet the needs of automobile companies to reduce the development cycle and budget. Moreover, CFD requires a high level of experience from development engineers, and in some cases, invalid calculation schemes often occur, and each high-fidelity CFD simulation may take a long time aultman2022evaluation ; biswas2019development ; jiang2021vehicle . As to wind tunnel tests, although the accuracy and reliability is quite good, it can only be used to test a small number of designs due to the need to create an equivalent scale model of the whole vehicle, which increases the time and cost consumed by the experiment. Deep learning-driven approaches can make it possible to speed up the design process and effectively evaluate aerodynamic designs by utilizing existing datasets to explore the aerodynamic performance of different vehicle designs baque2018geodesic ; remelli2020meshsdf ; rios2019scalability ; rios2021point2ffd ; song2023surrogate ; trinh20243d ; benjamin2025systematic .
Owing to intensive research in the field of computer vision, a variety of powerful deep learning techniques designed for processing three-dimensional (3D) models have been developed atzmon2018point ; guo2021pct ; li2018pointcnn ; qi2017pointnet ; qi2017pointnet++ ; wang2019dynamic . These studies have introduced advanced functionalities such as object recognition, classification, and the automatic construction of new models. By leveraging deep learning as an innovative approach, researchers are now able to evaluate aerodynamic performance more efficiently. Through the use of existing model data, they can construct datasets and carry out training, thereby significantly reducing the time and cost typically associated with traditional CFD simulations remelli2020meshsdf ; ribeiro2020deepcfd ; rios2019scalability ; song2023surrogate .
However, current technology has to face a large number of challenging problems in processing three-dimensional (3D) models with higher complexity. For example, the deployment of large deep learning models is limited by the lack of data sets and training resources. At the same time, there is still a significant gap between the prediction accuracy and generalization ability of smaller networks, which makes it difficult to fully meet the actual needs of current production processes jacob2021deep ; tychola2024deep ; vinodkumar2024deep ; wu2024towards .
A significant amount of research has been conducted on basic two-dimensional (2D) models such as airfoils jun2020application ; thuerey2020deep ; wu2020deep , and several mature deep learning methods have been summarized. Among these, DeepCFD, proposed by Ribeiro ribeiro2020deepcfd , utilizes deep learning models like Convolutional Neural Network (CNN) to quickly simulate the velocity and pressure fields of 2D steady-state laminar flow. The deep learning architecture of CFD Net, proposed by Obiols obiols2020cfdnet , leverages extensive model data to predict fluid characteristics rapidly under various boundary conditions and flow field environments, thereby reducing the computation time required by traditional CFD methods. Although these studies can enhance computational speed by several orders of magnitude compared to traditional CFD methods while maintaining lower error rates, they primarily focus on 2D steady laminar flow. There is a scarcity of research related to real-life applications involving complex models. In an effort to advance research, Song proposed proxy modeling of a 3D car model using a novel 2D representation to predict the drag coefficient song2023surrogate . Jacob introduced a different approach by employing a deep learning model to predict the aerodynamic drag coefficient of arbitrary vehicle shapes, addressing a limitation of traditional methods jacob2021deep . This method utilizes an improved U-Net architecture combined with a Signature Distance Field (SDF) to represent the input geometry, enabling accurate predictions of both drag coefficient and velocity fields after training. However, it was noted that the computational cost of augmenting the geometry is high. Given the current computational resources, it may not be practical for the network to learn to recognize the same geometric features from different directions in large-scale datasets. Meanwhile, Charles R. Qi proposed the PointNet architecture, a deep learning network specifically designed for processing 3D point cloud data qi2017pointnet . Its key innovation involves directly handling unordered point sets by processing each point independently and using symmetric functions (e.g., max-pooling) to extract global features, ensuring the network remains insensitive to the arrangement of the input points while remaining invariant to rigid-body transformations. This is significantly different from traditional convolutional neural networks, which rely on structured input data (e.g., images), and enables the extraction of model-specific parameters by analyzing 3D point cloud data. Dynamic Graph Convolutional Neural Network (DGCNN), proposed by Wang wang2019dynamic , is a specialized deep learning framework for 3D point cloud data processing. Its core idea is to characterize the point cloud data by graph structure and introduce dynamic graph convolutional operations to capture local and global geometric features. Unlike traditional static graph methods, DGCNN dynamically updates neighborhood relationships at each layer, allowing the model to flexibly learn complex geometric relationships within point cloud data at multiple scales. The fundamental EdgeConv operation further enhances modeling of local geometric information given the point cloud, effectively improving the accuracy and efficiency of feature extraction. DGCNN has been widely used in classification, segmentation and object recognition tasks involving 3D point cloud data, demonstrating superior performance in handling irregular geometric data. Subsequently, Rios investigated the use of 3D point cloud self-encoders to extract local geometric features in vehicle design optimization rios2021point2ffd . It was found that 3D point cloud self-encoders provide more efficient shape generation and improve the complementary degrees of freedom of the model, leading to improved optimization of aerodynamic performance.
Recently, Elrefaie proposed RegDGCNN, an extended version for regression tasks based on DGCNN elrefaie2024drivaernet . Unlike DGCNN, for the classification task, RegDGCNN employs a regression loss function (e.g., mean square error loss) to optimize the model for scenarios requiring continuous value predictions. This adaptation allows RegDGCNN to excel in applications such as aerodynamic parameter prediction, which demand precise numerical regression. By inheriting the dynamic graph convolution structure of DGCNN, RegDGCNN achieves both efficient feature extraction and accurate regression prediction performance when processing point cloud data with complex geometries, thus supporting a wide range of regression tasks. At its core, RegDGCNN processes point cloud data through EdgeConv. Although it performs well in extracting local geometric features, it is less effective in predicting continuous values. While this method shows strong capabilities in capturing local geometric characteristics, its limitation lies in the inability to effectively model the structural features of the entire shape, which may lead to inaccuracies in aerodynamic calculations. After all, aerodynamic coefficients are influenced not only by local morphology but also by the complex interactions between different components of the overall geometry. Further research by Elrefaie demonstrated that although the RegDGCNN architecture may achieve high prediction accuracy for identical vehicle configurations, its performance significantly declines when applied to models outside the training dataset elrefaie2024drivaernet++ . For entirely new models not present in the dataset, its prediction accuracy falls short of that achieved by CFD simulations, as the EdgeConv-based approach cannot fully leverage the available geometric information, resulting in suboptimal performance for complex shapes. This limitation highlights the urgent need for further advancements in the utilization of geometric data to improve predictive accuracy.
In view of the challenges revealed by the above research, this study proposes a rather novel method that combines the cloud-driven attention mechanism (CDA) with the correlation estimation module (CDE). This integrated approach aims to model the inherent morphological information in the disordered cloud structure in a more effective manner, thereby significantly improving the prediction accuracy and ability of various vehicle designs. This research has resulted in the development of a comprehensive system that can predict the aerodynamic drag coefficient of 3D STL vehicle models. Notably, the system can achieve a prediction accuracy close to that of CFD, while greatly reducing the time and resources generally required for traditional simulations. In addition, after optimization, the system can support the prediction of aerodynamic performance for a variety of vehicle designs, thereby further improving the generality and accuracy of the algorithm.
II DrivAerNet++ Dataset
Figure 1: Variation car body types include estateback, fastback, and notchback combined with open and closed wheels.
DrivAerNet++ is the largest and most comprehensive multi-modal 3D dataset available, specifically designed for data-driven aerodynamic analysis in automotive engineering elrefaie2024drivaernet++ . It includes 8,000 industry-standard automotive design models, each sample comes with a detailed 3D model (with up to 500,000 surface meshes) and provides extensive 3D flow fields and precise aerodynamic coefficients. This wealth of information enables researchers and engineers to predict vehicle aerodynamic drag metrics more accurately under various design conditions. With a 333% increase in data volume compared to the previous largest publicly available automotive dataset, DrivAerNet++ is the only open database that includes models of tires and chassis, providing a robust foundation for advanced aerodynamic calculations.
Moreover, DrivAerNet++ offers significant advantages over traditional aerodynamic datasets, such as ShapeNet and other open-source datasets song2023surrogate . Traditional datasets typically include only a limited number of vehicle designs and lack the geometric detail necessary for accurate aerodynamic simulations—especially in the modeling of components like wheels and chassis, which are critical for precise prediction. For instance, vehicle models in ShapeNet are often simplified into single-structure representations, failing to capture the complexity and diversity found in real-world vehicle designs. These simplified models tend to exhibit poor generalization performance when applied to different vehicle types and may even introduce bias.
In addition, Fig. 1 presents several example vehicle models from the DrivAerNet++ dataset. These models cover a wide range of vehicle types and design styles, showcasing both high fidelity in geometric detail and broad coverage across different car categories.
II.1 CFD Simulation Setup
The study utilizes the DrivAerNet dataset generated through computational fluid dynamics (CFD) simulations implemented in the open-source OpenFOAM® framework elrefaie2024drivaernet ; elrefaie2024drivaernet++ ; greenshields2022openfoam . A pressure-velocity coupled steady-state solver ($simpleFoam$) optimized for incompressible turbulent flow analysis was employed. The computational domain, constructed with 1:1 scale representation of the DrivAer fastback vehicle configuration, features dimensions of $5.0L\times 3.5L\times 2.5L$ (where $L$ denotes vehicle length), incorporating symmetry boundary conditions along the $y$-normal plane to enhance computational efficiency.
II.2 Turbulence Modeling
Turbulence closure was achieved through Menter’s modified $k$-$\omega$ SST model, which employs blending functions to transition smoothly between near-wall $k$-$\omega$ formulation and free-stream $k$-$\epsilon$ methodology menter2003ten . This approach enhances flow separation prediction accuracy under adverse pressure gradients. Governing equations encompass turbulent kinetic energy transport, specific dissipation rate, and eddy viscosity models. Boundary conditions include: uniform inflow velocity (30 m/s, corresponding to Reynolds number $\text{Re}=9.39\times 10^{6}$ based on vehicle length), zero-gradient velocity with fixed static pressure at outlet (configured with anti-backflow treatment), no-slip conditions on vehicle surfaces, rotating wall boundaries for wheels, and slip conditions for domain top/sides. Near-wall resolution was maintained using Spalding-law-based $nutUSpaldingWallFunction$ to ensure $Y^{+}\approx 1$ launder1983numerical .
II.3 Grid Generation
Hex-dominant meshes generated via $SnappyHexMesh$ contained 15 prism layers with initial height 0.1 mm and expansion ratio 1.2. Four-tier local refinement zones were implemented:
  • Vehicle proximity ($\Delta x=5$ mm)
  • Wake core ($\Delta x=10$ mm)
  • Lateral flow ($\Delta x=20$ mm)
  • Far-field background ($\Delta x=50$ mm)
Grid independence was confirmed with $<1\%$ drag coefficient variation between baseline (8M cells, 88 CPUh) and refined (16M cells, 205 CPUh) meshes.
II.4 Validation Results
Validation against experimental data for DrivAer fastback configuration showed:
  • Drag coefficient discrepancies: 2.81% (coarse) vs 0.81% (fine)
  • Surface pressure correlation $R^{2}>0.95$ with wind tunnel measurements
  • Wake velocity profile RMS error $<3\%$ compared to PIV data
confirming numerical model reliability heft2012experimental ; wieser2014experimental .
The aerodynamic drag coefficient $C_{d}$ is mathematically defined as:
$C_{d}=\frac{F_{d}}{\frac{1}{2}\rho u_{\infty}^{2}A_{\text{ref}}}$ (1)
where $F_{d}$ denotes the cumulative resistance force acting on the automotive structure, $\rho$ specifies the ambient air mass density, $u_{\infty}$ characterizes the undisturbed flow velocity, and $A_{\text{ref}}$ indicates the standardized projection area. The resultant force constitutes two distinct physical mechanisms: aerodynamic pressure differential effects (commonly termed shape resistance) and viscous shear forces induced by boundary layer development.
II.5 Feature Extraction
In this study, an Alpha Shape-based contour feature extraction algorithm is employed to capture the streamlined characteristics of 3D vehicle models edelsbrunner1994three . This approach leverages multi-layer sectional analysis along the vehicle’s longitudinal axis to extract high-density aerodynamic representations. A series of evenly spaced slicing planes are generated along the driving direction of the vehicle. The geometric intersections between these planes and the vehicle body yield cross-sectional contours, details are shown in the Fig. 2 and Fig. 3.
(a) Raw model (with reduced point cloud size)
(b) Model after contour extraction
Figure 2: DrivAer model along the yz-plane
(a) Raw model (with reduced point cloud size)
(b) Model after contour extraction
Figure 3: DrivAer model along the xz-plane
For each cross-section, the Alpha Shape algorithm is applied to automatically identify edge feature points. A curvature-sensitive and adaptive sampling strategy is then utilized: in regions where the curvature radius falls below a predefined threshold—such as the windshield transition zone and rear wing edges—the sampling density is increased by a factor of three compared to flatter areas. Ultimately, 1,000 feature points are extracted per section to form a sparse point cloud that preserves the aerodynamic characteristics of the vehicle.
This method can control a variety of variables such as the number of cross-sectional layers, sampling density, and curvature threshold in this specific way of parameterization. Throughout the implementation process, it is an important prerequisite to retain the overall vehicle shape, and also to ensure that the point density in key feature areas is higher than that of the traditional Poisson sampling method. The experimental results show that this method can significantly simplify the modeling of irrelevant structural details such as interior parts, radiators, and engines. It is important to note that this simplification process does not affect the prediction accuracy, and ultimately provides a more efficient solution for simulating the drag of the entire vehicle in industrial applications.
III DrivAer Transformer (DAT)
Figure 4: A regression framework for predicting drag coefficient. The model processes input mesh points with a 3D automotive mesh model based on a conventional point cloud feature extraction framework qi2017pointnet ; qi2017pointnet++ . The network employs L edge convolution along with a correlative attention fusion module to aggregate point set features and characterize boundary features through a dynamically learned graph network you2022roland . The associative attention module utilizes a multi-layer perceptron (MLP) with shared weights to determine edge features, constructing similarity and evaluation matrices. These auto-correlation matrices are dynamically weighted within the dynamic graph space feature extraction module, and the final MLP maps the aggregated 1D feature vectors into drag coefficient.
In the developmental research of geometric deep learning in recent years, the application of this method to solve hydrodynamic problems of complex geometries has a good research prospectkashefi2022physics ; pfaff2020learning ; rios2019scalability ; rios2021point2ffd ; rios2019efficiency ; sanchez2020learning . In this study, a series of deep learning modules for processing 3D point cloud data are designed and implemented aiming to efficiently predict the aerodynamic drag coefficient $C_{d}$ of a car through techniques such as geometric feature extraction, dynamic convolution and self-attention mechanism. The model is capable of extracting complex geometrical structures and topological information from 3D point clouds, which greatly enhances the accuracy of aerodynamic parameter prediction. The detailed architecture is illustrated in Fig. 4 and the original diagram is showed in appendix.
III.1 A regression prediction framework based on dynamic graph networks
Aerodynamic drag is strongly influenced by the morphology of the vehicle, so the prediction of aerodynamic drag needs to capture the local correlation of the point cloud data and the complex correlation effect between the global point cloud features. However, 3D point cloud data is a kind of disordered 3D data, in order to maintain the consistency of the learning results while using convolutional neural network for feature extraction of disordered point sets, the maximum pooling is used to aggregate the disordered point set features to form the global point set features. Moreover, for the local vehicle structure, the local geometric structure is utilized by applying a convolution-like operation on the edges connecting neighboring point pairs by constructing bowtie graph values, which enables the model to construct the local point cloud into a graph neural network thus capturing the key aerodynamic parameters of the fluid around the object more accurately. In this paper, on the basis of RegDGCNN, we add the point cloud structure correlation-driven attention (CDA) and correlation estimation module (CDE) to further model the morphological information of the disordered point cloud structure, which more accurately exploits the graph neural network features of the local tie points. Meanwhile, the spatial coding capability of PointNet extracted from the original point cloud features and the local relational inference of DGCNN are utilized to achieve the prediction of continuous values of aerodynamic parameters, and the feature bridges extracted by correlation attention perform more prominently in terms of translational invariance and nonlocality . After multiple layers of correlation edge convolution blocks, the constructed local bowtie graph changes in order to dynamically adapt to the changing point set feature space. In this study, the initialization graph $G$ and a point cloud feature set $X=\{x_{1},\ldots,x_{n}\}\in\mathbb{R}^{F}$ are defined. The graph feature $\gamma\in\mathbb{R}^{C\times P}$, where $C$ represents the number of channel dimensions, and $P$ represents the number of points.
III.2 Correlation-Driven Attention(CDA)
Since edge convolution (EdgeConv) can effectively model the geometric features of locally disordered point cloud structures when processing point cloud data, it is especially good at capturing semantic information in local regions, however, it still has some limitations in extracting global semantic structure features, especially for the application of predicting drag coefficient in aerodynamic tasks. This limitation is mainly reflected in the fact that edge convolution mainly focuses on the relationship between local feature points and ignores the global semantic structure of the overall point cloud, while the prediction of air drag coefficient often relies on the overall geometrical shape and the interactions between the parts. Therefore, the dynamic graph model constructed based on edge convolution is limited in its ability to extract global information and cannot fully utilize the global structural features of the point cloud, thus affecting the accurate prediction of aerodynamic performance.
To address this problem, this paper proposes an innovative attention mechanism based on point cloud feature invariance, which aims to further enhance the extraction of global semantic information through matrix transposition operations. In this method, we introduce a learned attention score at the level of channel features to dynamically regulate the importance of features in different channels. Specifically, by weighting the important channels, local features are made to be highlighted or suppressed in specific channels, which can capture the key parts of the global structure of the point cloud more effectively. In this way, the model can not only continue to play the advantage of edge convolution in local feature extraction, but also make up for its shortcomings in global information extraction through the channel attention mechanism.
In addition, the correlation-based attention mechanism proposed in this paper calculates the similarity between local graph features and using the similarity matrix formed between the transposed graph and the original input graph, the model’s ability to capture global geometric semantic information is significantly enhanced. The specific process is to calculate the similarity between the transposed graph features and the input features to generate a similarity matrix, which in turn effectively represents the global relationship between points. In order to further improve the expressive ability of the model, a new activation method is designed in this paper to enhance the expressive ability of the model by adjusting the nonlinear characteristics through the learnable parameters $p_{1}$, $p_{2}$, and $\beta$. This flexibility allows the model to better adapt to different input features. The following formulas illustrate the approach, assuming the AconC operator input $x\in\mathbb{R}^{1\times C\times 1}$ and learnable parameter $p_{1}\in\mathbb{R}^{1\times C\times 1}$:
$dpx$ $=(p_{1}-p_{2}\times x)$ (2)
$AconC$ $=dpx\times\sigma(\beta\times dpx)+p_{2}\times x$ (3)
$G^{q}$ $=AconC(\tau(\text{Conv}(\gamma^{T}))^{T})$ (4)
$G^{k}$ $=AconC(\tau(\text{Conv}(\gamma^{T})))$ (5)
$S$ $=(G^{q})^{T}\times G^{k}$ (6)
where $S$ represents the similarity matrix, $\tau$ denotes batch normalization, $\sigma$ is the Sigmoid activation function, and $\gamma^{T}$ represents the transposed input point cloud features.
III.3 Correlation-Driven Estimator(CDE)
On the basis of having obtained the similarity matrix $S$, this paper further proposes a novel method for expanding local similarities with specific properties into a global structural feature that represents a specific category. The idea is to treat the features carried by each channel as a representation of the corresponding global information. In detail, the function of local similarity information is to reflect the characteristics of the local geometric structure of the point cloud. However, in order to better capture the global geometric patterns and their distribution in the unordered point cloud, it is necessary to derive a representation of the global from the local similarity. This whole process actually involves processing the features contained in each channel, and also reallocating the channel weights to reflect the importance of the global features.
The method introduced for this purpose is based on the similarity matrix $S$. It re-calculates the weight of each channel by making full use of the affinity between channels. One of the most critical operations is to perform the $\arg\max$ operation on the similarity matrix $S$ for a specific dimension of the channel, so that the most prominent similarity features can be extracted from the similarity matrix $S$, that is, the maximum value within the channel dimension. The role of these extracted maxima is to indicate the most important information in each channel that the model needs to pay attention to, and it is of great help in identifying the parts that play a key role in the overall global structure.
Next, the features obtained through the argmax operation are normalized using the Softmax activation function, which enhances the influence of these important features in the channel dimension and ensures the smoothness and stability of the weights assigned.The nonlinear activation of Softmax effectively improves the differentiation of the features, which allows the model to focus on the most representative part of the global geometry. The above process can be expressed by the following equation. The above process can be expressed by the following equation:
The above process can be expressed as:
$A=\sigma\left(\arg\max(S)-S\right)$ (7)
where $\arg\max$ denotes the operation that extracts the maximum values along the channel dimension from the similarity matrix, obtaining the most critical information. $\sigma$ represents the Softmax non-linear activation function. This method enhances the model’s ability by extending local similarity to global features, and also improves the model’s ability to capture the overall structural characteristics of unordered point clouds by reallocating channel weights. When dealing with complex point cloud data, the introduced channel affinity mechanism can adaptively emphasize the channels that dominate the global features according to the actual situation, thereby enhancing the model’s performance. This performance is particularly prominent in task scenarios involving global geometric semantic information, and the mechanism also has a significant effect on improving the prediction accuracy of complex scenarios such as aerodynamics.
III.4 Dynamic learnable integration
While edge convolution provides local spatial features of point clouds, the correlation attention module introduced in this study offers global semantic features. To enhance the robustness of the fusion between these features, we adopt variable learnable parameters $\alpha$ with specific initialization to weight the fusion of local and global features. The fusion process is expressed as:
$G^{v}=AconC(\gamma)$ (8)
$G^{o}=\alpha\cdot A\cdot G^{v}+\gamma$ (9)
where $G^{o}$ represents the output features extracted by the regression prediction network.
III.5 Training and Evaluation of Models
The training of the DAT model adopts Mean Squared Error (MSE) as the primary loss function, combined with the Adam optimization algorithm kingma2014adam for parameter optimization. During the training process, strategies such as learning rate adjustment and L2 Regularization (or Weight Decay) are employed to prevent model overfitting.
The model’s performance is evaluated using metrics including $R^{2}$ score (Coefficient of Determination), Mean Absolute Percentage Error (MAE), and Mean Squared Error (MSE), which comprehensively reflect the model’s predictive capability and robustness. To accurately assess the model’s performance, the following three evaluation metrics are used: MSE, MAE, and $R^{2}$, as shown in Equations (11) to (13).
MSE $=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-y_{i}^{\prime})^{2}$ (10)
MAE $=\frac{1}{n}\sum_{i=1}^{n}\left|\frac{y_{i}-y_{i}^{\prime}}{y_{i}}\right|$ (11)
$R^{2}$ $=1-\frac{\sum_{i=1}^{n}(y_{i}-y_{i}^{\prime})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}$ (12)
where: $y_{i}$ represents the ground truth value, and $y_{i}^{\prime}$ is the predicted value.
The closer MSE and MAE are to 0, and the closer $R^{2}$ is to 1, the higher the prediction accuracy of the model.
IV Validation and Analysis
IV.1 Deficiencies in validation sets
For the DrivAerNet++ dataset, this study first allocates 70% of the dataset for training and 15% for validation and testing respectively. However, this dataset has some problems in the validation process:
  1. 1.
    The shape variations in DrivAerNet++ are small, but the number of samples is large, which can easily affect the training results, and more models need to be added for validation.
  2. 2.
    The mesh resolutions in the DrivAerNet++ dataset are almost the same, which may lead to aerodynamic prediction of models at other resolutions to have issues.
  3. 3.
    The DrivAerNet++ dataset was obtained by morphing using four car models as the main template, which does not take into account the fact that in reality, different car companies have their own design styles, which may lead to bias in the prediction.
IV.2 Additions to the DrivAerNet++ validation set
China’s rapid economic development has led to a gradual shift in customer demand for cars with large space designs, with SUVs (Sports Utility Vehicles), ORVs (Off-Road Vehicles) and MPVs (Multi-Purpose Vehicles) in particular in demand du2018insights ; helveston2015will . The current DrivAerNet++ dataset does not adequately cover these three key vehicle categories and is not fully consistent with the trend of the Chinese automotive market. In order to more accurately reflect the market demand and improve the applicability of the model in the Chinese market, this study plans to increase the data of SUVs and MPVs in the training set to make up for the lack of model diversity in the current validation set, thus enhancing the generalization ability of the model. Since the DrivAerNet++ dataset is derived from the base model DrivAer, only individual models are added to the PCV (Passenger Car Vehicle) part of this study for validation, and the rest of the main models are combined by sampling the DrivAerNet++ dataset. Some of the selected SUV (Sport Utility Vehicle), ORV (Off-Road Vehicle), MPV (Multi-Purpose Vehicle)and PCV models used are shown in Fig. 5 and 6.
Figure 5: Models Sample Showcase: A. SUV; B. ORV; C. MPV
Figure 6: PCV Models Sample Showcase: A. Self-driving cars; B. Tesla Model-S; C. CAERI Aero Model
Based on the results of The First Automotive CFD Prediction Workshop (AutoCFD1), a discussion of the Reynolds Averaged Navier-Stokes (RANS) method for multiple turbulence models, Wall Model Large Eddy Simulation (WMLES), and the RANS-LES Hybrid Method (HRLM) revealed that the HRLM and RANS methods differed significantly from the experimentally derived predictions of air resistance ashton2022overview . Therefore, in order to ensure the accuracy of the validation set, it was decided to use wind tunnel test data instead of CFD simulation data in this study, and some of the models were chosen from car models provided by relevant car companies cooperating with this research lab. In this study, a real vehicle model was scanned using 3D scanning technology and reverse modeled using Materialise Magics industrial software to generate a workable 3D solid model. This approach accurately captures the details and geometric features of the vehicle shape and ensures that the model is highly consistent with the actual vehicle. Next, these 3D solid models were imported into ANSA software for pre-processing, which included mesh delineation, geometric restoration and setting boundary conditions. Finally, the validation set stl file for testing can be obtained.
This study aims to construct an additional validation dataset in a situation that enhances the realism of the validation set and tests the broad applicability of the proposed modeling framework. Construct an additional validation dataset using wind tunnel experiments as a reliable aerodynamic test method that can provide accurate air resistance coefficient information and effectively improve the prediction performance and reliability of the model. Due to the high cost of large-scale wind tunnel experiments, the validation dataset used in this study was constructed using a combination of existing wind tunnel test data and supplementary experimental data. The wind tunnel equipment used is located in the Hubei Provincial Key Laboratory of Modern Automotive Parts Technology at Wuhan University of Technology, and its specific configuration and experimental conditions are shown in the Fig. 7.
Figure 7: Wind Tunnel
In the wind tunnel test, the vehicle model was fixed to ensure that the body attitude and position were unchanged to ensure the reproducibility of the results and the accuracy of the data. The car tires were kept immobile during the experiment to avoid airflow interference affecting the conclusions. At the same time, the seams of the model were sanded to make the surface smooth and reduce the deviation caused by air resistance. These design and control measures enhance the reliability of the test results. The specific operation steps and equipment arrangement are shown in Fig. 8.
Figure 8: Wind Tunnel Testing
From the Fig. 9(a), it is concluded that the drag coefficient $C_{d}$ of the DrivAerNet++ validation set is uniformly distributed and ranges from 0.2 to 0.4. The distribution of the extended validation set in different drag coefficient intervals is shown in Fig. 9(b) and Fig. 10. It can be seen that there is a large proportion of samples with 0.30-0.34 as the main region. This result is consistent with the aerodynamic characteristics of mainstream SUV and MPV vehicles in the market, and also shows that the supplemental dataset improves the model’s ability to generalize to various vehicle types ekman2020assessment ; wang2019effect ; windsor2014real . This extension not only makes up for the insufficient samples of SUV and MPV models in the DrivAerNet++ dataset, but also provides a more targeted support force for the application of the DAT architecture in the Chinese automotive market.
(a) DrivAerNet++ Validation Set
(b) New Validation Set
Figure 9: Comparison of Validation Sets
Figure 10: Normal distribution of New Validation Set
(a) Correlation of Predicted vs. Ground Truth $C_{d}$
(b) DrivAerNet++ Test Set’s Error Distribution
Figure 11: Error Analysis on the DrivAerNet++ Test Set
(a) Correlation of Predicted vs. Ground Truth $C_{d}$
(b) Error Distribution on the New Test Set
Figure 12: Error Analysis on the New Test Set
V Results and discussion
V.1 Results
The test results of PointNet qi2017pointnet , GCNN kipf2016semi , RegDGCNN elrefaie2024drivaernet and DAT under the validation set of DrivAerNet++ are shown in Tab. 1, and it can be seen that the MSE and MAE of the four algorithms don’t differ much from each other, and that DAT is improved compared to the other three algorithms in terms of $R^{2}$, and that the training time and the inference time don’t differ much from RegDGCNN.
Table 1: The test results of PointNet, GCNN, RegDGCNN and DAT.
Model MSE MAE Max AE $R^{2}$ Training Time
PointNet 0.000194 0.0115 0.0199 0.751 13.7 hrs
GCNN 0.000201 0.0132 0.0223 0.774 51.3 hrs
RegDGCNN 0.000186 0.0112 0.0176 0.833 14.7 hrs
DAT 0.000178 0.0105 0.0166 0.871 15.5 hrs
After evaluating the validation set, it was found that the DAT architecture was more adaptable to a variety of models and that the self-attention vaswani2017attention component significantly improved the accuracy of the model’s predictions for the dataset. Comparing the distribution of RegDGCNN and DAT in different prediction error intervals in Fig. 11, it can be seen that the DAT architecture has an advantage in the low error range. The prediction data of RegDGCNN model has a more even error distribution, while the prediction data of DAT model is more located in a smaller error range, most of the data is within 4%.
Under the new validation set obtained by using wind tunnel experimental data, the data obtained by DAT and RegDGCNN architectures are shown in Fig. 12.
(a) Predicted $C_{d}$ on E_S_WW_WM Set
(c) Predicted $C_{d}$ on F_D_WM_WW Set
(e) Predicted $C_{d}$ on F_S_WWS_WM Set
(g) Predicted $C_{d}$ on N_S_WW_WM Set
(b) Predicted $C_{d}$ on E_S_WWC_WM Set
(d) Predicted $C_{d}$ on F_S_WWC_WM Set
(f) Predicted $C_{d}$ on N_S_WWC_WM Set
(h) Predicted $C_{d}$ on N_S_WWS_WM Set
Figure 13: Comparison of predicted $C_{d}$ and ground truth on eight test sets from DrivAerNet++
(a) PCV Error Distribution
(b) SUV Error Distribution
(c) ORV Error Distribution
(d) MPV Error Distribution
Figure 14: Comparison of PCV, SUV, ORV, and MPV Error Distributions
The results of the study show that the DAT architecture has better generalization to brand new models outside the training set. This distributional feature shows that the DAT architecture is able to provide more accurate predictions on the expanded dataset. Fig. 13 presents the complete test dataset from DrivAerNet++, showing the predicted aerodynamic drag coefficient ($C_{d}$) against the ground truth values. The results show that the average prediction accuracy of the DAT model significantly exceeds that of the RegDGCNN model regardless of the kind of model’s validation set, validating the potential and advantages of the DAT architecture in terms of design and accuracy improvement.
Since the newly constructed dataset contains numerous car models, the data predicted by different types of cars were classified in this study to obtain Fig. 14. As shown in the Fig. 14, the DAT architecture outperforms the RegDGCNN architecture in the validation of several car models, especially in the MPV and ORV models, which suggests that the Self-Attention module has a better result in predicting the models that have a large deviation from the model in the training set, which is exactly the difficult point where machine learning could not be applied to real production in the past. In Tab. 2, this study compares the error of wind resistance predicted by RegDGCNN, DAT architecture with the real data obtained from wind tunnel experiments, and it can be seen that DAT can reduced the error from 4%-5% to 2%-3% for RegDGCNN. Also their high accuracy performance has met the requirements of most automobile manufacturers for the prediction accuracy of drag coefficient of automobiles.
Table 2: $C_{d}$ Range and Error Rates for Different Vehicle Models.
Model $C_{d}$ Range RegDGCNN DAT
PCV 0.26 - 0.34 4.650% 2.553%
SUV 0.28 - 0.35 5.099% 3.374%
ORV 0.32 - 0.40 4.805% 3.916%
MPV 0.31 - 0.38 5.428% 3.025%
Relying on the excellent self-attention module, the DAT architecture shows strong adaptive ability and high prediction accuracy on the newly constructed dataset, which indicates that it has a wide range of applications in the field of automobile aerodynamics prediction.
V.2 Discussion
After analyzing the two architectures RegDGCNN and DAT from a variety of car model datasets, the research results show that DAT has a better generalization to untrained datasets and the computational accuracy meets the needs of the existing automobile development, and this feature is an important step for the application of machine learning methods in real production. However, our team encountered the following problems when validating a variety of models:
  1. 1.
    As smart driving technology began to be gradually applied to family cars, many models began to be retrofitted with smart driving devices, such as LIDAR, ultrasonic radar and vision camera components on the original model, and a comparison of the retrofitted car model is shown in Fig. 15. Such additional components cannot be well adapted to the change by either RegDGCNN or DAT architectures, and the specific parameters are shown in Tab. 3. It can be clearly seen that the car retrofitted with intelligent driving devices is then predicted with a significant deviation, which is far beyond what is acceptable in the car development process.
    (a) Basic Model
    (b) Enhanced Model
    Figure 15: Error Distribution Across Different Vehicle Types
    Table 3: Error Rate Comparison between Basic and Enhanced Models.
    Model DAT RegDGCNN
    Basic Model 0.298 (3.84%) 0.284 (8.68%)
    Enhanced Model 0.304 (10.85%) 0.298 (12.61%)
  2. 2.
    For some minor structural changes, such as adjusting the angle of the rear spoiler, changing the shape of the mirrors, etc., and adding built-in radiator parts. The machine learning method may not be able to accurately predict the change trend after the change, for example, the change of the angle of the rear spoiler will make the drag coefficient of the whole car produce a change of decreasing and then increasingcheng2019experimental ; das2017cfd . After comparing the data through deep learning calculations, it is found that such subtle trend changes are likely to be covered by the bias generated by each prediction to the extent that a generalized pattern cannot be derived. This study hypothesizes that this is a phenomenon that may be due to insufficient training samples.
VI Conclusion
In this paper, we obtained data on the aerodynamic performance of automobiles through numerical simulation and high-fidelity CFD simulation, and constructed a large-scale multimodal dataset, DrivAerNet++, containing different design parameters and aerodynamic coefficients. Eventually, this study established a prediction model DAT based on geometric deep learning and verified the differences of different models without using test set samples. The algorithm accomplishes innovations in the following two points: 1. Utilizing the negative feedback regulation mechanism in the field of automatic control, the idea of error-correcting feedback structure to comprehensively capture the local features of the point cloud; 2. Utilizing the attention module based on the affinity of the channel to help avoid possible redundancy in feature mapping. The conclusions of the study are as follows:
  1. 1.
    Using the DrivAerNet++ dataset and the geometric deep learning model, this study is able to efficiently and accurately predict the drag coefficient of a car. The model’s error is controlled within 4% under both the initial validation set and the supplementary validation set, showing high prediction accuracy and greater generalizability to other models relative to the RegDGCNN and DGCNN architectures. Moreover, the modified algorithm can directly process 3D mesh data, and the model does not require additional image rendering or generation of symbolic distance function (SDF), which simplifies the preprocessing steps and improves the computational efficiency and practicality of the model.
  2. 2.
    By extending the validation dataset with more data from SUV, MPV and other models, this study enhances the generalization ability of the DAT model. In these extended datasets, the DAT architecture outperforms the traditional RegDGCNN model, showing greater adaptability and higher prediction accuracy. For different air drag coefficient $C_{d}$ value intervals, the prediction error of the DAT model is reduced. This indicates that the model works better in handling new data and improves in the prediction of aerodynamic performance of complex 3D automobile shapes. The results after validation using wind tunnel experiments to obtain accurate data show that the DAT model not only meets the current automobile manufacturing industry’s requirements for prediction accuracy, but also demonstrates promising applications on large-scale multimodal datasets.
  3. 3.
    Although the DAT model demonstrated excellent performance in experiments, the study also identified some potential problems in the training and optimization process, such as limited computational resources, high data processing complexity, and high cost of wind tunnel testing. Future research should focus on streamlining the data management and loading strategies, improving the efficiency of computational resource usage, and exploring more cost-effective experimental validation methods to promote the popularization of the DAT architecture in practical engineering applications. Through continuous optimization and extension, the model is expected to gain wide and far-reaching influence in automotive aerodynamic performance prediction and other related fields.
Acknowledgment
This work was supported by the National Natural Science Foundation of China (Grant No. 52472416).